Write the slope-intercept equation of the line l that goes through (12, 3) and has a slope of 1/3.

Answer by Alan3354(69443)

Write the slope-intercept equation of the line l that goes through (12, 3) and has a slope of 1/3.
--------------
These are easier than they look. Use the standary slope-intercept form (y = mx + b) with the one point.
y - y1 = m*(x - x1)
y-3 = (1/3)*(x-12)
Multiply by 3 to get rid of fractions
3y-9 = x-12
x - 3y = 3
That's all.